OpenStudy (anonymous):

Combine like terms.B) 2y^2-8y+9-12y^2-8y+3

OpenStudy (zpupster):

just arrange like terms 2y^2-12y^2-8y-8y+9+3 then add them together watch your signs!!

OpenStudy (anonymous):

2y^2-12y^2= -8y-8y= 9+3 =

OpenStudy (zpupster):

yes thats it